### v1

# Free Your Hands: Lightweight Relightable Turntable Capture Pipeline

Abstract.
Novel view synthesis (NVS) from multiple captured photos of an object is a widely studied problem. Achieving high quality typically requires dense sampling of input views, which can lead to frustrating and tedious manual labor. Manually positioning cameras to maintain an optimal desired distribution can be difficult for humans, and if a good distribution is found, it is not easy to replicate. Additionally, the captured data can suffer from motion blur and defocus due to human error. In this paper, we present a lightweight object capture pipeline to reduce the manual workload and standardize the acquisition setup. We use a consumer turntable to carry the object and a tripod to hold the camera. As the turntable rotates, we automatically capture dense samples from various views and lighting conditions; we can repeat this for several camera positions. This way, we can easily capture hundreds of valid images in several minutes without hands-on effort. However, in the object reference frame, the light conditions vary; this is harmful to a standard NVS method like 3D Gaussian splatting (3DGS) which assumes fixed lighting. We design a neural radiance representation conditioned on light rotations, which addresses this issue and allows relightability as an additional benefit. We demonstrate our pipeline using 3DGS as the underlying framework, achieving competitive quality compared to previous methods with exhaustive acquisition and showcasing its potential for relighting and harmonization tasks.
Corresponding authors. Email: csjyang@njust.edu.cn
Corresponding authors. Email: beibei.wang@nju.edu.cn
submissionid: 826
1. Introduction
Creating realistic 3D digital assets from real-world objects is a long-standing challenge of computer graphics, with applications to e-commerce, entertainment, digital heritage and more. A typical pipeline captures multiple views of an object under fixed lighting and transforms the captured images into a 3D digital asset, which allows novel view synthesis (NVS) or in some cases, relighting. Neural radiance fields (NeRF) (Mildenhall et al., 2021) or 3D Gaussian splatting (3DGS) (Kerbl et al., 2023) can be used for this purpose, but high-quality photograph capture still requires careful manual labor and time, asthese methods require fairly dense views (typically hundreds). Handheld camera videos are faster to capture but frequently suffer from motion blur and defocus.
Images can also be captured with specialized devices (Ye et al., 2024; Ma et al., 2021; Kang et al., 2023); this can reduce the required human labor and precision, but these devices are often not easily accessible, limiting their applicability. In the case of hand-held cameras, various lighting setups can be employed, including unknown environment lighting (Zhang et al., 2020; Wang et al., 2021), or a point light that is collocated with the camera (Bi et al., 2020c, b, a), or separate from the camera (Gao et al., 2020; Bi et al., 2024a). In all cases, the camera is held by a human hand. The constraints on a reasonable sampling of viewpoints may be obvious to researchers, but manually positioning cameras to maintain the desired distribution can be difficult for average users. Additionally, the quality of captured data can suffer from issues like motion blur and defocus, as shown in Fig. 2. Several works propose utilizing sparse views and cross-scene feed-forward inference techniques (Chen et al., 2025; Charatan et al., 2024). However, the reconstructed quality necessarily degrades without dense samples.
Figure 2. Illustrations of some common problems in the ordinary hand-hold capturing pipeline. The final NVS quality can suffer from the imperfectly sampled distribution and blurry images.
In this paper, we present a lightweight object capture pipeline designed to reduce manual workload, standardize the acquisition process and make it repeatable, while also obtaining some relightability as an additional benefit. We use a consumer turntable to carry the object and a tripod to support the camera. As the turntable rotates, we automatically capture dense samples from various views and under different lighting conditions (i.e., rotating environment light as observed from the object’s frame of reference). This way, we can easily capture hundreds of valid images in several minutes without hands-on effort, while minimizing issues with blur and defocus. The captured images can be used in existing NVS frameworks (NeRF or 3DGS). However, the varying light conditions break the assumptions of these techniques, and are harmful to the reconstruction quality. Therefore, we design a radiance representation conditioned on light rotations to bridge the gap.
We first discuss the insights behind our turntable design from a theoretical perspective. Next, we demonstrate our pipeline using 3DGS as the underlying framework, achieving competitive quality compared to previous methods with exhaustive acquisition, while also showcasing potential for relighting and harmonization tasks. To summarize, our main contributions include:
  • a novel lightweight capture pipeline that uses a turntable and tripod to minimize manual workload and standardize the acquisition process,
  • a rotation-conditioned radiance field representation tailored for our capture pipeline and allowing relightability in the form of lighting rotation, and
  • a dataset consisting of glossy and furry objects captured and reconstructed under different lighting configurations with our pipeline, available for downstream research.
2. Related work
Figure 3. Intuitively, varying view directions and light rotations can be thought of as a 2-dimensional space of variation. The typical NVS problem (left) can be seen as sampling a single 1D “column” in this space. In our formulation (middle) and its planar projection (right), we use a turntable setup to easily obtain samples along the “diagonal lines” in this two-variable sample space. We capture multiple circles as shown in the right sub-figure, and each circle makes up an orbit of cameras, leading to a continuous and uniform sampling of viewpoints. We fit a manifold to these samples, enabling us to render novel views as well as light rotations; as a consequence, we can finally solve the NVS problem in a specific 1D slice of this manifold.
Novel view synthesis
Novel view synthesis (NVS) aims to generate new images from view directions that were not originally observed. Extensive research on 3D representations has been proposed to enable realistic novel view rendering. Notably, NeRF (Mildenhall et al., 2021), 3DGS (Kerbl et al., 2023) and their follow-ups have garnered significant attention due to their powerful representation capabilities.
NeRF models radiance using integrals over a ray passing through a volume, and addresses view-dependent radiance using multilayer perceptrons (MLPs) conditioned on the view ray direction; this is also true in most follow-up methods (Chen et al., 2022; Müller et al., 2022). Enhanced anti-aliasing techniques (Barron et al., 2021, 2022, 2023; Zhang et al., 2020) and improved reflectance modeling (Verbin et al., 2022; Attal et al., 2023) have further refined the quality and performance of NeRF-based representations. 3DGS (Kerbl et al., 2023) employs anisotropic Gaussians to represent scenes, allowing for great adaptivity to actual geometric content and enabling real-time, highly detailed renderings. The view dependence of radiance is represented using spherical harmonics, which is even more limited than MLP-based view-dependence.
Inverse rendering and relighting
Based on these NVS frameworks, some inverse rendering approaches (Zhang et al., 2021; Yao et al., 2022; Liu et al., 2023; Bi et al., 2020a; Boss et al., 2021; Jin et al., 2023; Liang et al., 2024) allow for the relighting with novel environment lights by decomposing the light, material and geometry of the target object. Gao et al. (2020) leverage a learned neural texture on a rough proxy geometry to achieve relighting of objects. Mullia et al. (2024) propose a novel representation that combines explicit geometry with a neural feature grid and an MLP decoder, achieving high-fidelity rendering and relighting with good flexibility and integration. NRHints (Zeng et al., 2023) maintains an implicit neural representation with both SDF and NeRF and predicts radiance with shadow and highlight hints, achieving high-quality relighting. With 3DGS as the framework, some work (Bi et al., 2024b; Gao et al., 2023) also relight objects with analytical appearance approximation or neural appearances. These approaches decompose the illumination and materials from the object, supporting arbitrary novel environments or point lights with full relightability, but at the cost of introducing limited material models and often baking lighting features into material properties. Compared to these methods, we opt for a simpler way without explicitly modeling the illumination, achieving partial relighting with rotated or linearly combined environments.
Camera pose calibration
These above techniques rely on posed images to densely reconstruct 3D scenes, with the quality of the input images significantly influencing the final rendering results. Structure from Motion (SfM) methods (Hartley and Zisserman, 2003; Schonberger and Frahm, 2016; Mur-Artal et al., 2015; Taketomi et al., 2017), particularly COLMAP (Schönberger and Frahm, 2016), are commonly employed to calibrate camera poses and provide initialization for point-based techniques. Although the quality of input data is typically assumed to be sufficient for COLMAP to succeed, this assumption is frequently invalid. Consequently, standardizing the data acquisition process and minimizing human error is beneficial to ensure robust and reliable performance in NVS applications.
3. Our Method
To create a 3D asset from a real-world object for novel view synthesis, a common approach is to capture hundreds of images by moving the camera around the object under a static light condition and then reconstruct the radiance field with NeRF or 3DGS. The entire pipeline can be formulated as follows:
(1) $L=\mathcal{G}(\mathbf{V_{\mathrm{s}}},I_{0}),$
where $L$ represents the radiance field, $\mathcal{G}$ denotes a reconstruction operator, which can be either NeRF or 3DGS. $\mathbf{V_{s}}$ is a set of camera views parameterized by the camera intrinsics and extrinsics. $I_{0}$ is an arbitrary static lighting (illumination) setting (e.g., unknown environment) under which the images are captured.
To ensure high-quality reconstruction, it is necessary to have a dense set of samples, typically hundreds of camera views, leading to manual labor and high possibility of human error. Can we free our hands while maintaining high-quality reconstruction? To answer this challenge, we introduce a novel capture pipeline, presented below.
3.1. Capturing with a turntable
In addition to moving the camera, we can also move or rotate the object itself. A simple yet effective solution to this issue is to use a turntable to hold the object, allowing it to rotate automatically. We can position the camera on a tripod in a small number of locations, to easily capture hundreds of images.
Suppose that the camera is placed at $M$ different locations. For each camera location, a sequence of images corresponding to a full turntable rotation forms a circle of $N$ camera positions in the object frame. Unlike the typical capture pipeline, the images are captured under different light conditions when thinking in the object frame of reference: the same lighting environment is rotated differently with respect to the object. This leads to our conditional radiance field reconstruction formulation:
(2) $L$ $=\mathcal{G}(\mathbf{V},\mathbf{I}),$
(3) $\{\mathbf{V}\}$ =$=\{\mathbf{V}_{i}|i=0,1,...,M\},$
(4) $\{\mathbf{I}\}$ =$=\{\mathbf{I}_{j}|j=0,1,...,N\},$
where $\mathbf{V}$ consists of samples from $M$ circles captured by placing the camera around the turntable at different elevation and azimuth angles. Here each $\mathbf{V}_{i}$ is a set of views from one circle, and $\mathbf{I}$ is a set of lighting rotations rather than a static lighting ($I_{0}$ in Eqn. (1)). We choose $N$ rotations ($\mathbf{I}_{j}$) from each circle, resulting in a total of $M\times N$ samples.
With this setup, we can easily capture hundreds of images with minimum human labor. The remaining question is how to reconstruct the conditional radiance field with these captured images.
3.2. Radiance field reconstruction with varying lighting conditions
The original formulation for radiance field reconstruction, as presented in Eqn.  (1), essentially operates as a one-variable regression problem, predicting radiance values from different viewpoints. In contrast, our conditional radiance field reconstruction approach incorporates a rotating light environment (in object frame), making it a regression problem involving two variables. The key insight behind our formulation is to construct a field conditioned on the two-variable sample space and interpret the NVS problem as a one-dimensional slice of this field, as illustrated in Fig. 3.
We use 3DGS as our underlying 3D representation, though NeRF formulations could be used with small modifications as well. In standard 3DGS, radiance is represented using spherical harmonics (SH) basis functions. While SH vectors are a compact way to represent angular distributions, there is no option to make them conditional on varying light conditions. Instead, we replace per-Gaussian SH representation with a neural network conditioned on light rotations (in addition to view directions) and build the manifold in a neural latent space.
Specifically, each Gaussian has an appearance latent feature vector representing the observed color from given views and light rotations. This appearance latent vector is interpreted by an MLP to decode the conditional radiance field. The color per Gaussian is computed as:
(5) $c=\mathrm{MLP}(x|v,\theta),$
where $x$ is the latent vector stored in each Gaussian point, $v$ is the view direction (a normalized 3D direction), and $\theta$ is the azimuthal rotation angle. Specifically, the MLP is 128-channel with 2 hidden layers, using a level-2 frequency encoding for the input view directions and light rotations. The latent vectors are 8-dimensional. We evaluate the network before splatting the Gaussians, and the final pixel colors are computed using standard 3DGS differentiable rasterization.
3.3. Applications of our capture pipeline
Our capture pipeline can be applied to various scenarios. Here, we outline several of these applications.
NVS and relighting with the neural radiance representation
One obvious application of our pipeline is for NVS, as stated in the previous section, by computing the radiance for each Gaussian using its latent vectors and performing the alpha-blending. Additionally, since our neural presentation is defined by all different light rotations, it naturally facilitates relighting under all rotating light conditions.
NVS with a distilled SH representation
Besides using the neural representation directly, it can be distilled into other representations (e.g., SHs) by fixing a light rotation angle. The distilled representation allows novel view synthesis, while being compatible with existing 3DGS-based applications.
Relighting under novel light conditions
Since our neural representation encodes the radiance field with multiple light conditions, we can relight the object by linearly combining different light rotations (e.g., front-lit and back-lit ones), achieving a novel light condition. Furthermore, if the original environment lighting is known, it is possible to approximately fit a target lighting using the known rotations, in order to achieve even more complex relighting tasks.
4. Implementation details
Turntable capture dataset
We have constructed a dataset of turntable captures under various indoor lighting conditions. The dataset consists of 17 real-world objects with diffuse, furry, or glossy appearances. Each object is sampled using $M=8$ turntable videos featuring different camera elevations and azimuths, and we select $N=60$ frames from each video, resulting in a total of 480 images for each training set. Another 60 frames from a video at a novel view is used as the testing set. A typical capture process takes 3 minutes.
Our pipeline is flexible and allows for the use of controlled photographic studio lighting or additional camera samples. We offer a version of our dataset captured under a studio light box that delivers controllable soft lighting. Such extensions can enhance the result quality and repeatability, while maintaining the efficiency and simplicity of our pipeline. We describe the studio box variant in Fig. 4 and validate it in Sec. 5. We also provide another variation with 20 extra views captured under static lighting, and we discuss it in our supplementary.
Figure 4. The illustration of some alternative versions of our datasets. The users can easily extend the setup of our pipeline with their desired extensions without obviously increasing the workload and time cost. We showcase and discuss the extra sparse static captures setup in our supplementary.
Capturing details and data preparation
We show a snapshot of our capturing environment in Fig 1. We use the 48 mm video camera from an iPhone 15 Pro Max and crop the images to get a resolution of 1440$\times$1440 pixels. During capturing, we position the camera approximately every $0.25\pi$ radians around the object to create 8 video sequences, while uniformly varying elevation angles. For the studio box case, we only sample viewpoints from the front side of the box, since the object is blocked from other sides.
After capturing, we compute optical flow to identify the start and end points of each rotation cycle (though of course we could also manually stop and restart the videos). We randomly select 60 frames from each cycle. The turntable rotates at a fixed angular velocity, so we divide $2\pi$ by the total number of frames to determine the rotation angle at each selected timestamp. All images are masked and the background is automatically removed using SAM2 (Ravi et al., 2024). We then conduct calibration using COLMAP (Schönberger and Frahm, 2016).
Optimization
During optimization, we back-propagate the gradient to the MLP weights, feature vectors and Gaussian parameters. The network weights and the latent vectors are jointly updated and optimized to fit the appearance under the rotating light conditions with $L_{1}$ loss function and the structural similarity index (SSIM) (Wang et al., 2004).
Table 1. The rendering quality of our neural radiance representation and 3DGS with novel views and light rotations on our captured datasets (both default and studio box). The results are evaluated in PSNR ()($\uparrow$)( ↑ ), SSIM ()($\uparrow$)( ↑ ) and LPIPS ()($\downarrow$)( ↓ ). The best/second-best results are colored in red/orange. The soft light from the studio box generally provides higher qualities, while the capturing workload stays the same. Our neural radiance representation outperforms 3DGS on both variants of datasets.
Scene Default dataset (3 minutes) (Fig.8) Studio box dataset (3 minutes) (Fig.10)
Ours 3DGS Ours 3DGS
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Alpaca 33.7039 0.9894 0.0169 29.6081 0.9806 0.0323 34.0134 0.9896 0.0184 31.7581 0.9830 0.0344
Cat 30.9368 0.9835 0.0109 26.0663 0.9555 0.0356 40.8780 0.9938 0.0046 35.4237 0.9818 0.0133
China 34.2610 0.9897 0.0090 30.7809 0.9844 0.0146 39.6296 0.9956 0.0052 36.5625 0.9933 0.0084
Controller 30.4726 0.9761 0.0209 25.2222 0.9659 0.0345 35.4382 0.9920 0.0089 33.0897 0.9856 0.0180
Dinosaur 36.0936 0.9928 0.0065 32.5989 0.9883 0.0123 41.9424 0.9967 0.0045 38.4615 0.9934 0.0097
Dolphin 39.0878 0.9961 0.0048 34.1935 0.9924 0.0086 42.5473 0.9972 0.0041 40.2936 0.9953 0.0075
Jar 29.3733 0.9553 0.0484 23.6761 0.9101 0.1196 33.1215 0.9772 0.0193 28.8963 0.9415 0.0798
Monza 36.5302 0.9926 0.0054 29.6788 0.9844 0.0104 41.3471 0.9964 0.0034 38.6690 0.9926 0.0077
Panda 29.1212 0.9780 0.0288 25.7316 0.9628 0.0525 36.3338 0.9783 0.0112 32.5720 0.9614 0.0266
Pearl 38.0879 0.9917 0.0070 30.4644 0.9834 0.0140 41.4716 0.9960 0.0045 37.4667 0.9916 0.0105
Penguin 28.9595 0.9698 0.0290 23.9183 0.9453 0.0598 35.2508 0.9792 0.0194 33.0846 0.9571 0.0456
Pine 25.6992 0.9451 0.0421 21.7535 0.8820 0.0881 32.1439 0.9724 0.0184 28.0977 0.9354 0.0580
Rabbit 31.6668 0.9760 0.0200 29.4303 0.9592 0.0413 34.1400 0.9842 0.0191 30.5215 0.9648 0.0413
Rider 28.6987 0.9486 0.0345 25.3948 0.9275 0.0535 34.9932 0.9816 0.0190 32.7340 0.9640 0.0378
RiderSmall 37.5932 0.9924 0.0053 31.0377 0.9830 0.0115 41.9490 0.9961 0.0055 38.5647 0.9917 0.0105
Sparrow 29.7554 0.9720 0.0309 25.6735 0.9593 0.0531 31.9495 0.9585 0.0132 31.1654 0.9526 0.0260
Tuan 31.5752 0.9827 0.0234 23.9322 0.9693 0.0439 34.6477 0.8956 0.0087 25.7017 0.8694 0.0277
Average 32.4968 0.9785 0.0202 27.4832 0.9609 0.0403 37.1645 0.9812 0.0110 33.7096 0.9679 0.0272
Table 2. The comparison of NVS qualities on synthetic datasets. All results are evaluated in PSNR ()($\uparrow$)( ↑ ) and LPIPS ()($\downarrow$)( ↓ ), and best/second-best results are marked in red/orange. Note that in this comparison, the testing set is without any light rotations. Our method produces higher-quality results from training images with rotating lights than 3DGS does with common dataset setups (100 static images). Even with fairly sufficient ideally sampled static samples, our results are still competitive with them.
Scene Ours 3DGS 3DGS
(600 rotating) (100 static) (660 static)
PSNR LPIPS PSNR LPIPS PSNR LPIPS
Armadillo 35.760 0.025 35.949 0.036 36.588 0.035
Ficus 34.227 0.007 34.001 0.008 35.207 0.007
Flowers 31.920 0.020 30.237 0.037 31.288 0.033
Lego 33.507 0.026 32.137 0.045 32.957 0.041
Average 33.853 0.020 33.081 0.032 34.010 0.029
5. Results
5.1. Experiment setup
We validate our pipeline on our captured and synthetic datasets. We run all experiments on an Ubuntu 22.04 LTS distribution powered by Windows Subsystem Linux 2. All renderings are done with $800\times 800$ resolution, and training our pipeline typically takes about 15 minutes on an RTX4090 GPU. We evaluate the peak signal-to-noise ratio (PSNR), SSIM, and perceptual similarity (LPIPS) (Zhang et al., 2018) values to compare both the pixel-wise accuracy and global degree of realism.
5.2. Datasets and baseline
We run experiments on our captured read-world datasets and synthesized datasets. As stated in Sec. 4, we provide two versions of our dataset with different capturing setups: (a) using the turntable only (called the default dataset) and (b) using an additional studio box for soft lighting (called the studio box dataset). We evaluate the NVS and relighting qualities on both datasets in this section. Note that in both datasets, each image is with different views and light rotation angles. We build synthetic data with a simulated turntable only to validate the NVS quality. In the synthetic dataset, all training images are with light rotations but testing images, and we generate ideally sampled fully static images for 3DGS to compare with.
We use 3DGS (Kerbl et al., 2023) as our baseline for comparison. Since other advanced approaches (Meng et al., 2024; Liu et al., 2024; Yu et al., 2024) that enhance 3DGS are orthogonal to our pipeline, we did not perform comparisons with these methods.
Figure 5. The NVS result with SHs distilled from our trained model. By a simple distillation process, our conditional radiance representation can be exported into a static radiance field at a specific light rotation angle for simple NVS and easy cooperation with other applications.
Figure 6. Relighting the objects by linear combinations of light rotations with RGB weights. Top: blue and purple light combined. Bottom: yellow and green light combined.
Figure 7. The effect of the number of circles ($M$) in our pipeline. Given only one fixed camera position, we only observe one “diagonal line” in the view-light space, resulting in severe over-fitting. In practice, we choose $M=8$ for our real datasets.
5.3. Quality validation
To validate the effectiveness of our pipeline, we compare the rendering quality between our conditional radiance representation and 3DGS on default and studio box datasets, where quantitative results are provided in Table 1 and the visual comparisons are provided in Figs. 8 and 10. Generally, the studio box datasets give a higher rendering quality, since the softened light makes it easier for both our method and 3DGS to represent the radiance. Our method outperforms 3DGS on both datasets, showing $\sim$32 dB on the default dataset and $\sim$37 dB on the studio box dataset.
Default dataset
In Fig. 8, we provide the visual comparison between our method and 3DGS across three scenes selected from the default datasets. These scenes include typical complex fur and reflections. By comparison, we find that our method has a much higher quality (about +5 dB) consistently. In particular, 3DGS exhibits an overly dark appearance at grazing angles and obvious blurriness on the furry surface as it cannot handle rotating light conditions. Although it can reconstruct good geometries, the radiance (especially the shadow effects) is incorrectly predicted. In contrast, our method produces results that closely match the reference, thanks to the conditional radiance representation.
Studio box dataset
We further compare the rendering quality between our method and 3DGS on four scenes from the studio box datasets, which offer high-quality photographic lighting with less lighting variation. These scenes are chosen to cover several typical effects, including high specular, detailed fur, etc. Our method still surpasses 3DGS in all these scenarios, showing about 5 dB higher PSNR on average. While the quality of 3DGS has also been improved compared to the ones from the default datasets, it still suffers from missing specular highlights (top left) and over-blurred appearance (two scenes at the bottom), while our renderings preserve these details and better match the ground truth.
We provide results from synthetic datasets using our capturing pipeline but test with full static images (without any light rotations) to compare the NVS quality. In this comparison, 3DGS uses ideally sampled fully static training images. We provide the rendering comparisons in Fig. 9 and quantitative results in Table 2. Our neural radiance representation achieves higher-quality renderings than that from 3DGS with a common setup (100 static images) and is still competitive with results of 3DGS from abundantly sampled static views (660 static images). Note that in practice, even for 600 images with rotating lights, we still need no more than 4 minutes to capture them, while capturing 100 high-quality static images already costs more than 3 minutes of difficult manual labor. In addition, our NVS quality can be further improved with another variant of our datasets. We describe and validate this in our supplementary.
5.4. More applications
Relighting with the neural representation
Since our method reconstructs a two-dimensional manifold, considering both the light rotations and different views, it naturally supports relighting with the same light but with different rotations. In Fig. 11, we showcase the rendered result under different light rotations. We use the Rabbit scene from the default dataset as an example. With varying lighting conditions, our rendered results exhibit overall reasonable lighting outcomes. Since it is difficult to achieve rotating light with a fixed camera in practice, we don’t provide ground truths for comparison in this figure.
NVS with distilled SH representation
After selecting a lighting angle, our neural representation can be easily distilled into an SH representation. This way, it can fit existing 3DGS-based applications. In Fig. 5, we show the NVS results rendered with the distilled SHs under three different lighting rotations. The distilled SHs achieve close qualities to the results from our neural radiance representation, producing faithful NVS compared to the ground truth at the specified rotation angles. Note that we obtain all the results from only one training. A typical distillation process can finish within 3 minutes.
Relighting under novel light conditions
Our neural representation encodes the radiance field with multiple light conditions, which allows us to linearly combine different light rotations (e.g., front-lit and back-lit ones), achieving a fused light condition. In Fig. 6, we demonstrate the relighting results by combining the front-lit and the back-lit scenarios, with different lighting colors. The way, our method enables the creation of rich and diverse assets.
5.5. Ablation study
In our capture pipeline, we need to position the camera on a tripod in several locations, where this location number $M$ is a crucial factor. To validate its influence, we provide an ablation study in Fig. 7 by setting different values of $M$ with the default setup. For fairness, we keep the same total number of captures across the different setups (1, 4, and 8 locations). By comparison, increasing the number of locations improves the rendering quality. The main reason behind this phenomenon is that sparse locations lead to uneven sampling on the 2D manifold, which can lead to overfitting during optimization and hurting the rendering quality.
5.6. Discussion and limitations
More flexible relighting
Our capture pipeline enables relighting with different light rotations or a linear combination of various lighting conditions. However, our method still cannot fully or flexibly support relighting in entirely new environments. We will address this in future work.
Data processing
Our capture pipeline relies on COLMAP (Schönberger and Frahm, 2016) to calibrate our camera poses. However, the light condition changes during the capturing, which raises the difficulty of calibrating. While the current background removal can alleviate this issue, it might occasionally fail due to the capability of the large model. Therefore, either a COLMAP-free calibration or a more accurate background removal approach will further improve our capture quality.
Alternative NVS frameworks
We choose the basic 3DGS framework to validate the effectiveness of our capturing pipeline. However, there are still many advanced approaches that provide even higher NVS qualities. Their work is orthogonal to ours and our pipeline is flexible enough to fit into most of such advanced approaches.
6. Conclusion
In this paper, we have presented Free Your Hands, a lightweight object-capturing pipeline to reduce manual workload, standardize the acquisition process, and ensure repeatability. The proposed capture pipeline consists of a simple setup: a consumer turntable to hold the target object and a tripod to hold the camera. As the turntable rotates, we can easily capture hundreds of valid images in several minutes without hands-on effort. Then, we design a neural radiance representation conditioned on light rotations tailored for the captured images. Our capture pipeline can be integrated into both NeRF or 3DGS-based frameworks. We have demonstrated the effectiveness of our pipeline in the 3DGS-based framework across various applications, including NVS and relighting under different light rotations or combined lighting conditions, showing competitive quality in NVS and reasonable relighting effects.
There are still many potential future research directions. One promising avenue is to expand the neural radiance representation into a fully relightable manifold. Additionally, improving the NVS and relighting quality on some difficult types of objects, such as reflective or transparent ones, is also an interesting and challenging direction.

### v3

@@
 Additionally, improving the NVS and relighting quality on some difficult types of objects, such as reflective or transparent ones, is also an interesting and challenging direction.
+Supplementary Materials:
+Free Your Hands: Lightweight Relightable Turntable Capture Pipeline
+Table 3.
+Rendering qualities with novel view and light rotation from our neural radiance representation on different datasets.
+The results are evaluated in PSNR ()($\uparrow$)( ↑ ), SSIM ()($\uparrow$)( ↑ ) and LPIPS ()($\downarrow$)( ↓ ).
+The best/second-best/third-best results are colored in red/yellow.
+Our neural radiance representation outperforms 3DGS on both variants of datasets.
+Scene Default dataset (3 minutes) (See the main paper) +20 static dataset (3.7 minutes) (Fig.12)
+Ours 3DGS Ours 3DGS
+PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
+Alpaca 33.7039 0.9894 0.0169 29.6081 0.9806 0.0323 34.2296 0.9906 0.0161 29.1272 0.9802 0.0329
+Cat 30.9368 0.9835 0.0109 26.0663 0.9555 0.0356 33.1386 0.9828 0.0114 26.0425 0.9535 0.0415
+China 34.2610 0.9897 0.0090 30.7809 0.9844 0.0146 35.3399 0.9903 0.0095 30.3217 0.9835 0.0153
+Controller 30.4726 0.9761 0.0209 25.2222 0.9659 0.0345 30.7255 0.9767 0.0206 24.5290 0.9648 0.0356
+Dinosaur 36.0936 0.9928 0.0065 32.5989 0.9883 0.0123 36.3225 0.9930 0.0062 33.5739 0.9889 0.0125
+Dolphin 39.0878 0.9961 0.0048 34.1935 0.9924 0.0086 39.1236 0.9959 0.0050 33.6032 0.9920 0.0091
+Jar 29.3733 0.9553 0.0484 23.6761 0.9101 0.1196 29.8483 0.9484 0.0596 23.7533 0.9012 0.1390
+Monza 36.5302 0.9926 0.0054 29.6788 0.9844 0.0104 36.5283 0.9928 0.0050 29.4366 0.9841 0.0101
+Panda 29.1212 0.9780 0.0288 25.7316 0.9628 0.0525 28.6421 0.9778 0.0288 26.4826 0.9660 0.0517
+Pearl 38.0879 0.9917 0.0070 30.4644 0.9834 0.0140 37.5966 0.9921 0.0066 30.5545 0.9839 0.0143
+Penguin 28.9595 0.9698 0.0290 23.9183 0.9453 0.0598 29.5938 0.9699 0.0295 23.9765 0.9467 0.0600
+Pine 25.6992 0.9451 0.0421 21.7535 0.8820 0.0881 29.2246 0.9504 0.0375 21.8756 0.8822 0.0922
+Rabbit 31.6668 0.9760 0.0200 29.4303 0.9592 0.0413 32.1203 0.9823 0.0198 29.8564 0.9657 0.0409
+Rider 28.6987 0.9486 0.0345 25.3948 0.9275 0.0535 28.5096 0.9498 0.0346 25.6186 0.9280 0.0538
+RiderSmall 37.5932 0.9924 0.0053 31.0377 0.9830 0.0115 37.2683 0.9922 0.0056 31.0191 0.9817 0.0128
+Sparrow 29.7554 0.9720 0.0309 25.6735 0.9593 0.0531 29.4498 0.9715 0.0316 25.7855 0.9596 0.0541
+Tuan 31.5752 0.9827 0.0234 23.9322 0.9693 0.0439 31.3717 0.9827 0.0239 23.6291 0.9688 0.0447
+Average 32.4968 0.9785 0.0202 27.4832 0.9609 0.0403 32.8843 0.9788 0.0206 27.5991 0.9606 0.0424
+Table 4.
+The NVS qualities with different numbers of extra sparse static captures as input.
+All resutls are evaluated in PSNR ()($\uparrow$)( ↑ ), SSIM ()($\uparrow$)( ↑ ), and LPIPS ()($\downarrow$)( ↓ ), and best/second-best/third-best results are marked in red/orange/yellow.
+Note that in this figure, the testing set is without any light rotations.
+By adding sparse captures with the targeted light condition, the NVS quality of our pipeline can be further improved, even surpassing 3DGS with the ideal exhaustive acquisition.
+However, we only need no more than 4 minutes in practice to obtain such a dataset, while it will take more than 20 minutes for a hand-held camera to finish the ordinary capturing.
+Scene Ours (600 rotating) Ours (+10 static) Ours (+30 static) 3DGS (100 static) 3DGS (660 static)
+PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
+Armadillo 35.760 0.968 0.025 36.812 0.971 0.023 38.175 0.974 0.021 35.949 0.963 0.036 36.588 0.965 0.035
+Ficus 34.227 0.984 0.007 34.896 0.986 0.006 35.900 0.988 0.005 34.001 0.984 0.008 35.207 0.988 0.007
+Flowers 31.920 0.976 0.020 32.434 0.978 0.019 32.657 0.979 0.019 30.237 0.963 0.037 31.288 0.970 0.033
+Lego 33.507 0.959 0.026 34.077 0.963 0.024 35.398 0.971 0.019 32.137 0.951 0.045 32.957 0.958 0.041
+Average 33.853 0.972 0.020 34.555 0.974 0.018 35.532 0.978 0.016 33.081 0.965 0.032 34.010 0.970 0.029
+Figure 12.
+The rendering results with novel views and light rotations on +20 static datasets.
+The best/second-best results are marked as bold/italic.
+Compared to the default setting, Our NVS quality can be further improved on such datasets, while 3DGS still has incorrect shadow effects.
+Figure 13.
+The ablation on the impact of extra sparse static captures.
+Our method can significantly benefit from sparse additional samples under the desired light conditions, and still beat the quality of 3DGS, despite providing as many uniformly sampled ideal captures under a static light condition for them.
+Note that in this figure, all references are with the static lighting (i.e., NVS without any light rotations).
+In the main paper, we have proposed Free Your Hands, a lightweight relightable object-capturing pipeline with a turntable, allowing for NVS and relighting.
+We have also introduced a novel formulation of NVS problems and provided two versions of datasets that contain 17 real-world objects.
+In this supplementary, we introduce and discuss another variant of our proposed dataset with 20 extra sparse static captures as input.
+Our pipeline captures images under rotating light conditions and achieves rendering with novel views and light rotations simultaneously.
+However, when focusing on NVS (without any light rotations), we can further improve the qualities by adding a few extra samples at the target fixed light rotation.
+Therefore, we also present a variant of our dataset with 20 extra images captured under the static environment light (called the +20 static dataset).
+Theoretically, more extra static captures bring higher NVS quality at the target light condition.
+In practice, to balance the time cost and quality, we capture 20 extra static images.
+We show the effect of these extra sparse static captures in Fig.
+12, and also provide the quantitative results on our +20 static datasets, compared to the default ones in Table 3.
+Generally, The +20 static datasets give similar qualities compared to default datasets on such testing sets since the extra captures will not help the rendering with novel light rotations, but only for the static NVS.
+To validate the improvement in NVS, we also provide results on synthetic datasets with static images as the testing set.
+We use +20 static for real-world datasets and use +10/30 static synthetic datasets to demonstrate the effectiveness of extra static captures and show the impact of the number of static captures.
+In practice, even for +30 static, we still only need about 1 minute extra work time (4 minutes in total), while the hand-held camera can only capture about 120 valid static images in the same amount of time.
+In Fig.
+13, we show the impact of extra static captures in the final NVS quality, and quantitative results are also provided in Table 4.
+Our NVS result from rotating captures is closer to the ground truth compared to a standard 3DGS capture (from 100 static images), and stays competitive with 3DGS using the ideal exhaustive acquisition (660 static images).
+With extra static captures, we can further improve our NVS quality, outperforming 3DGS even with 660 static captures.
+Additionally provided static captures significantly improve our NVS quality, since the network can focus more on the specific 1D sample space than the global manifold.
+In contrast, 3DGS only gets marginal improvement from extra captures.
+We only demonstrate this on synthetic scenes, since the ideal high-quality and reasonably sampled static captures for 3DGS are difficult to collect in practice.
+Our method can significantly benefit from sparse additional samples under the desired light condition without much extra workload and time cost.